OpenStudy (anonymous):

@jim_thompson5910 Graph the circle with center at (-4, -3), which also passes through the point (0, 3). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

You first need to find the distance from (-4,-3) to (0,3)

jimthompson5910 (jim_thompson5910):

so you would use the distance formula again

OpenStudy (anonymous):

can you set it please :)

jimthompson5910 (jim_thompson5910):

d = sqrt((x2-x1)^2+(y2-y1)^2)

jimthompson5910 (jim_thompson5910):

x1=-4 y1=-3 x2=0 y2=3

jimthompson5910 (jim_thompson5910):

that gives you d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((0-(-4))^2+(3-(-3))^2) d = sqrt((0+4)^2+(3+3)^2) d = sqrt((4)^2+(6)^2) d = ???

OpenStudy (anonymous):

16+36

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

52

OpenStudy (anonymous):

now what

jimthompson5910 (jim_thompson5910):

so the distance is \[\Large d = \sqrt{52}\]

jimthompson5910 (jim_thompson5910):

The radius is \[\Large r = \sqrt{52}\] So \[\Large r^2 = 52\]

OpenStudy (anonymous):

@jim_thompson5910 so the radius i have to go 52 directions?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{52} \approx 7.21110255\]

jimthompson5910 (jim_thompson5910):

So you roughly have to go 7.2 units in each direction starting at the center

OpenStudy (anonymous):

can you show me how you go 7.2 units?

jimthompson5910 (jim_thompson5910):

you'll have to go 7 units, then a bit more (not 8 units though) So it's between 7 and 8, but much closer to 7 (than 8)

OpenStudy (anonymous):

ok cool wheres my center?

OpenStudy (anonymous):

-4,3?

jimthompson5910 (jim_thompson5910):

(-4, -3)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

isnt it positive 3?

OpenStudy (anonymous):

oh nvm

jimthompson5910 (jim_thompson5910):

no it' negative

jimthompson5910 (jim_thompson5910):

it's*